# String graphs are quasi-isometric to planar graphs

### v1

@@ 4 Encasings
See Figure˜4 for an illustration of some $7$-caged $H$. We say that $\mathcal{B}$ is a $d$-cage of $(G,\mathcal{H)}$ if every $H\in\mathcal{H}$ is $d$-caged by $\mathcal{B}$.
Figure 4: An illustration of some $H\in\mathcal{H}$ (coloured red) that is 7-caged by $\mathcal{B}$ (coloured blue). For $u,v\in H$, the vertex sets $A_{0},\ldots,A_{7}$ (coloured green) are as in the definition of being 7-caged. As an example illustrating the end of the proof of Lemma˜4.2, we would have that $A_{0}=A_{u}$, $A_{1}\in\mathcal{F}_{u}$, $A_{2},A_{5}\in\mathcal{A}_{u,v}$, $\{A_{3},A_{4}\}=\mathcal{V}_{(X,Y)}$, $A_{6}\in\mathcal{F}_{v}$, and $A_{7}=A_{v}$.
The motivation behind the notion of a $d$-cage is that it will allow us to “cut up” connected sets $H$ in such a way that each part is in some sense still close via the encircling $\mathcal{B}$. Being able to do this is crucial to the inductive argument of Theorem˜2.5.
@@
Figure 5: A choice of $V^{1}_{(X,Y)},V^{2}_{(X,Y)}$ (coloured purple) as in the proof of Lemma˜4.2. Elements of $\mathcal{B}$ are coloured blue and elements of $\mathcal{H}$ are coloured red.
For a clean separation $(X,Y)$ of $G^{*}$, we let $H_{(X,Y)}$ be the vertices $v$ of $G$ contained in some vertex set of $\mathcal{F}\cap Y$ such that there exists some $H\in\mathcal{H}_{2}$ containing both $v$ and intersecting $\bigcup_{A\in X\cap Y}A$. If $(X,Y)$ is 1-clean, then we let $V_{(X,Y)}=H_{(X,Y)}\cup\bigcup_{R\in R_{(X,Y)}}R$. Note that in this case, $V_{(X,Y)}$ is a connected set of $G$ and $I_{\mathcal{H}}(V_{X,Y})$ has weak diameter at most $3\cdot 210=630$ in $\textbf{RIG}(G,\mathcal{H})$. Furthermore, if $v\in V_{(X,Y)}\backslash\left(\bigcup_{B\in\mathcal{B}}B\right)=V_{(X,Y)}\backslash\left(\bigcup_{H\in\mathcal{H}_{0}\cup\mathcal{H}_{1}}H\right)$, then there exists some connected set $H^{\prime}\subseteq H\in\mathcal{H}_{2}$ of $G$ that contains $v$, is contained in $V_{(X,Y)}$, and intersects $\bigcup_{H\in\mathcal{H}_{0}\cup\mathcal{H}_{1}}H$ (or more precisely, the unique vertex set of $X\cap Y$). Otherwise, if $(X,Y)$ is 2-clean with $X\cap Y=\{B^{1},B^{2}\}$, then we choose disjoint $V_{(X,Y)}^{1}$, $V_{(X,Y)}^{2}$ such that $B^{1}\subseteq V_{(X,Y)}^{1}$, $B^{2}\subseteq V_{(X,Y)}^{2}$, $V_{(X,Y)}^{1}\cup V_{(X,Y)}^{2}=H_{(X,Y)}\cup\bigcup_{R\in R_{(X,Y)}}R$, and both $V_{(X,Y)}^{1}$ and $V_{(X,Y)}^{2}$ are connected sets of $G$ (see Figure˜5 for an illustration). Then $I_{\mathcal{H}}(V_{(X,Y)}^{1}\cup V_{(X,Y)}^{2})$, and in particular both of $I_{\mathcal{H}}(V_{(X,Y)}^{1})$ and $I_{\mathcal{H}}(V_{(X,Y)}^{2})$ all have weak diameter at most $4\cdot 210=840$ in $\textbf{RIG}(G,\mathcal{H})$. Furthermore, if $v\in(V_{(X,Y)}^{1}\cup V_{(X,Y)}^{2})\backslash\left(\bigcup_{B\in\mathcal{B}}B\right)=(V_{(X,Y)}^{1}\cup V_{(X,Y)}^{2})\backslash\left(\bigcup_{H\in\mathcal{H}_{0}\cup\mathcal{H}_{1}}H\right)$, then there exists some connected set $H^{\prime}\subseteq H\in\mathcal{H}_{2}$ of $G$ that contains $v$, is contained in $V_{(X,Y)}^{1}\cup V_{(X,Y)}^{2}$, and intersects $B^{1}\cup B^{2}\subseteq\bigcup_{H\in\mathcal{H}_{0}\cup\mathcal{H}_{1}}H$.
@@
Consider some $u,v\in H$. Let $H_{u}$ be a subset of $H$ that contains $u$ and is maximal with respect to being a connected set of $G$ and being contained in $G[\bigcup_{B\in\mathcal{B}}B]$ (or equivalently $G[\bigcup_{B\in\mathcal{B}^{*}}B]$). Let $A_{u}$ be the union of all $B\in\mathcal{B}^{*}$ that intersect $H_{u}$. Let $\mathcal{F}_{u}$ be elements $F$ of $\mathcal{F}^{*}$ such that $H$ intersects both $F$ and $A_{u}$. We define $A_{v}$ and $\mathcal{F}_{v}$ similarly. Crucially, note that if $H$ intersects distinct $F_{1},F_{2}\in\mathcal{F}^{*}$, then $H$ intersects some $V\in\mathcal{B}$ incident to the face of $G(\mathcal{B}^{*})$ corresponding to $F_{1}$ that is also either contained in $\mathcal{V}_{(X,Y)}$ or adjacent to some element of $\mathcal{V}_{(X,Y)}$ in $G(\mathcal{B}^{*})$. Let $\mathcal{A}_{u,v}$ be the collection of unions of subcollections $\mathcal{B}^{\prime}\subseteq\mathcal{B}^{*}\backslash\mathcal{V}_{(X,Y)}$ such that some connected set $H^{*}\subseteq H$ of $G$ intersects each member of $\mathcal{B}^{\prime}$ and is contained in $G[\bigcup_{B\in\mathcal{B}^{\prime}}B]$, and some $B\in\mathcal{B^{\prime}}$ touches some $V\in\mathcal{V}_{(X,Y)}$. To see that $H$ is 7-caged by $\mathcal{B}^{*}$, it is now straight forward to choose the required sequence of vertex sets $A_{0},\ldots,A_{t}$ with $t\leqslant 7$ from the collection $\{A_{u}\}\cup\mathcal{F}_{u}\cup(\mathcal{A}_{u,v}\cup\mathcal{V}_{(X,Y)})\cup\mathcal{F}_{v}\cup\{A_{v}\}$ (here we take $A_{0}=A_{u}$, $A_{t}=A_{v}$, at most one $A_{i}$ is contained in $\mathcal{F}_{u}$ (this would be $A_{1}$ if so), at most one is contained in $\mathcal{F}_{v}$ (this would be $A_{t-1}$ if so), at most two are contained in $\mathcal{A}_{u,v}$, and at most two more are contained in $\mathcal{V}_{(X,Y)}$). For an illustrative example, see Figure˜4.
@@
See Figure˜6 for an illustrative example. We further say that some $H^{\prime}\in\mathcal{H}$ is $\mathcal{B}$-encroached by some $H\in\mathcal{H}$ if there is some set $F$ of vertices of $G$ contained in some face of $G[\bigcup_{B\in\mathcal{B}}B]$ and some vertex $v\in H^{\prime}\cap F$ such that $v$ is $\mathcal{B}$-encroached by $H$.
Figure 6: An encircling $\mathcal{B}$ (in blue) and some (red) $H\in\mathcal{H}$. The vertices of the facial vertex set $F$ that are encroached by $H$ are coloured orange.
@@ 5 String graphs
Figure 7: A fortification $G^{\prime}$ of planar graph $G$ (which is the subgraph on the black edges. The arrows orthogonal to the additional red edges indicate their sides.
@@
We remark that for the purposes of this paper, the last bullet could be weakened so that $H$ and $H^{\prime}$ just contain an end of $e$. However, this stronger version does not complicate the proof any further and we need this property for our applications to Riemannian surfaces [17].
@@
Clearly $\{v_{1},\ldots,v_{r}\}\subseteq R_{F,H}$. Let $E_{F,H}$ be the collection of such edges $v_{i}v_{i+1}$ (across each $v\in R_{F,H}$) that are not already an edge of $G$. For each such $v_{i}v_{i+1}\in E_{F,H}$, let $F_{v_{i}v_{i+1}}$ be the side of $v_{i}v_{i+1}$ drawn in the corresponding face of $G[\bigcup_{B\in\mathcal{B}}B]$ that contains the facial walk of $G$ between $u_{i}$ and $u_{i+1}$ that is contained in $U$ (note that the vertices of this facial walk are bounded $H$ in $G[\bigcup_{B\in\mathcal{B}}B]$ and the only two other vertices incident to this face are $v_{i},v_{i+1}$). Since $\mathcal{B}$ is a 4-surround of $(G,\mathcal{H})$, it follows that for every $H^{\prime}\in\mathcal{H}$ that contains a vertex incident to $F_{v_{i}v_{i+1}}$, the distance between $H$ and $H^{\prime}$ in $\textbf{RIG}(G,\mathcal{H}\backslash\mathcal{H}_{0})$ is at most $4$. This in particular implies that $F_{v_{i}v_{i+1}}$ will satisfy the 4th bullet in the definition of a $8$-fortification of $(G,\mathcal{H})$. This in particular implies that no vertex incident to $F_{v_{i}v_{i+1}}$ is incident to the outerface of $G$. Then observe that $R_{F,H}$ is a connected set in $G\cup E_{F,H}$. For an illustration of such a $R_{F,H},E_{F,H}$, see Figure˜8. Let $G^{*}$ be the planar graph obtained from adding these edges from the union of all such $E_{F,H}$. Let $\mathcal{R}$ be the collection of all such $R_{F,H}$. Then every element of $\mathcal{R}$ is a connected set of $G^{*}$.
Figure 8: The (orange) vertex set $R_{F,H}$ given the (blue) encasing $\mathcal{B}$, the (red) $H\in\mathcal{H}$, and $F\in\mathcal{F}$. The orange vertices are the vertices of $R_{F,H}^{\prime}$. The pink edges are $E_{F,H}$, and the arrows orthogonal to them indicate their sides.
@@
Let $G$ be a finite planar graph and $\mathcal{H}$ a spanning collection of connected vertex sets of $G$. Then there is a $(73936,80)$-impression $(G^{\prime},\mathcal{M})$ of $(G^{\prime},\mathcal{H})$, where $G^{\prime}$ is a fortification of $G$ and $\mathcal{H}$ 88-spans $G$.

### v2

@@ 4 Encasings
See Fig.˜4 for an illustration of some $7$-caged $H$. We say that $\mathcal{B}$ is a $d$-cage of $(G,\mathcal{H)}$ if every $H\in\mathcal{H}$ is $d$-caged by $\mathcal{B}$.
Рис. 4: An illustration of some $H\in\mathcal{H}$ (coloured red) that is 7-caged by $\mathcal{B}$ (coloured blue). For $u,v\in H$, the vertex sets $A_{0},\ldots,A_{7}$ (coloured green) are as in the definition of being 7-caged. As an example illustrating the end of the proof of Lemma˜4.2, we would have that $A_{0}=A_{u}$, $A_{1}\in\mathcal{F}_{u}$, $A_{2},A_{5}\in\mathcal{A}_{u,v}$, $\{A_{3},A_{4}\}=\mathcal{V}_{(X,Y)}$, $A_{6}\in\mathcal{F}_{v}$, and $A_{7}=A_{v}$.
The motivation behind the notion of a $d$-cage is that it will allow us to ‘‘cut up’’ connected sets $H$ in such a way that each part is in some sense still close via the encircling $\mathcal{B}$. Being able to do this is crucial to the inductive argument of Theorem˜2.5.
@@
Рис. 5: A choice of $V^{1}_{(X,Y)},V^{2}_{(X,Y)}$ (coloured purple) as in the proof of Lemma˜4.2. Elements of $\mathcal{B}$ are coloured blue and elements of $\mathcal{H}$ are coloured red.
For a clean separation $(X,Y)$ of $G^{*}$, we let $H_{(X,Y)}$ be the vertices $v$ of $G$ contained in some vertex set of $\mathcal{F}\cap Y$ such that there exists some $H\in\mathcal{H}_{2}$ containing both $v$ and intersecting $\bigcup_{A\in X\cap Y}A$. If $(X,Y)$ is 1-clean, then we let $V_{(X,Y)}=H_{(X,Y)}\cup\bigcup_{R\in R_{(X,Y)}}R$. Note that in this case, $V_{(X,Y)}$ is a connected set of $G$ and $I_{\mathcal{H}}(V_{X,Y})$ has weak diameter at most $3\cdot 210=630$ in $\textbf{RIG}(G,\mathcal{H})$. Furthermore, if $v\in V_{(X,Y)}\backslash\left(\bigcup_{B\in\mathcal{B}}B\right)=V_{(X,Y)}\backslash\left(\bigcup_{H\in\mathcal{H}_{0}\cup\mathcal{H}_{1}}H\right)$, then there exists some connected set $H^{\prime}\subseteq H\in\mathcal{H}_{2}$ of $G$ that contains $v$, is contained in $V_{(X,Y)}$, and intersects $\bigcup_{H\in\mathcal{H}_{0}\cup\mathcal{H}_{1}}H$ (or more precisely, the unique vertex set of $X\cap Y$). Otherwise, if $(X,Y)$ is 2-clean with $X\cap Y=\{B^{1},B^{2}\}$, then we choose disjoint $V_{(X,Y)}^{1}$, $V_{(X,Y)}^{2}$ such that $B^{1}\subseteq V_{(X,Y)}^{1}$, $B^{2}\subseteq V_{(X,Y)}^{2}$, $V_{(X,Y)}^{1}\cup V_{(X,Y)}^{2}=H_{(X,Y)}\cup\bigcup_{R\in R_{(X,Y)}}R$, and both $V_{(X,Y)}^{1}$ and $V_{(X,Y)}^{2}$ are connected sets of $G$ (see Fig.˜5 for an illustration). Then $I_{\mathcal{H}}(V_{(X,Y)}^{1}\cup V_{(X,Y)}^{2})$, and in particular both of $I_{\mathcal{H}}(V_{(X,Y)}^{1})$ and $I_{\mathcal{H}}(V_{(X,Y)}^{2})$ all have weak diameter at most $4\cdot 210=840$ in $\textbf{RIG}(G,\mathcal{H})$. Furthermore, if $v\in(V_{(X,Y)}^{1}\cup V_{(X,Y)}^{2})\backslash\left(\bigcup_{B\in\mathcal{B}}B\right)=(V_{(X,Y)}^{1}\cup V_{(X,Y)}^{2})\backslash\left(\bigcup_{H\in\mathcal{H}_{0}\cup\mathcal{H}_{1}}H\right)$, then there exists some connected set $H^{\prime}\subseteq H\in\mathcal{H}_{2}$ of $G$ that contains $v$, is contained in $V_{(X,Y)}^{1}\cup V_{(X,Y)}^{2}$, and intersects $B^{1}\cup B^{2}\subseteq\bigcup_{H\in\mathcal{H}_{0}\cup\mathcal{H}_{1}}H$.
@@
Consider some $u,v\in H$. Let $H_{u}$ be a subset of $H$ that contains $u$ and is maximal with respect to being a connected set of $G$ and being contained in $G[\bigcup_{B\in\mathcal{B}}B]$ (or equivalently $G[\bigcup_{B\in\mathcal{B}^{*}}B]$). Let $A_{u}$ be the union of all $B\in\mathcal{B}^{*}$ that intersect $H_{u}$. Let $\mathcal{F}_{u}$ be elements $F$ of $\mathcal{F}^{*}$ such that $H$ intersects both $F$ and $A_{u}$. We define $A_{v}$ and $\mathcal{F}_{v}$ similarly. Crucially, note that if $H$ intersects distinct $F_{1},F_{2}\in\mathcal{F}^{*}$, then $H$ intersects some $V\in\mathcal{B}$ incident to the face of $G(\mathcal{B}^{*})$ corresponding to $F_{1}$ that is also either contained in $\mathcal{V}_{(X,Y)}$ or adjacent to some element of $\mathcal{V}_{(X,Y)}$ in $G(\mathcal{B}^{*})$. Let $\mathcal{A}_{u,v}$ be the collection of unions of subcollections $\mathcal{B}^{\prime}\subseteq\mathcal{B}^{*}\backslash\mathcal{V}_{(X,Y)}$ such that some connected set $H^{*}\subseteq H$ of $G$ intersects each member of $\mathcal{B}^{\prime}$ and is contained in $G[\bigcup_{B\in\mathcal{B}^{\prime}}B]$, and some $B\in\mathcal{B^{\prime}}$ touches some $V\in\mathcal{V}_{(X,Y)}$. To see that $H$ is 7-caged by $\mathcal{B}^{*}$, it is now straight forward to choose the required sequence of vertex sets $A_{0},\ldots,A_{t}$ with $t\leqslant 7$ from the collection $\{A_{u}\}\cup\mathcal{F}_{u}\cup(\mathcal{A}_{u,v}\cup\mathcal{V}_{(X,Y)})\cup\mathcal{F}_{v}\cup\{A_{v}\}$ (here we take $A_{0}=A_{u}$, $A_{t}=A_{v}$, at most one $A_{i}$ is contained in $\mathcal{F}_{u}$ (this would be $A_{1}$ if so), at most one is contained in $\mathcal{F}_{v}$ (this would be $A_{t-1}$ if so), at most two are contained in $\mathcal{A}_{u,v}$, and at most two more are contained in $\mathcal{V}_{(X,Y)}$). For an illustrative example, see Fig.˜4.
@@
See Fig.˜6 for an illustrative example. We further say that some $H^{\prime}\in\mathcal{H}$ is $\mathcal{B}$-encroached by some $H\in\mathcal{H}$ if there is some set $F$ of vertices of $G$ contained in some face of $G[\bigcup_{B\in\mathcal{B}}B]$ and some vertex $v\in H^{\prime}\cap F$ such that $v$ is $\mathcal{B}$-encroached by $H$.
Рис. 6: An encircling $\mathcal{B}$ (in blue) and some (red) $H\in\mathcal{H}$. The vertices of the facial vertex set $F$ that are encroached by $H$ are coloured orange.
@@ 5 String graphs
Рис. 7: A fortification $G^{\prime}$ of planar graph $G$ (which is the subgraph on the black edges. The arrows orthogonal to the additional red edges indicate their sides.
@@
Clearly $\{v_{1},\ldots,v_{r}\}\subseteq R_{F,H}$. Let $E_{F,H}$ be the collection of such edges $v_{i}v_{i+1}$ (across each $v\in R_{F,H}$) that are not already an edge of $G$. For each such $v_{i}v_{i+1}\in E_{F,H}$, let $F_{v_{i}v_{i+1}}$ be the side of $v_{i}v_{i+1}$ drawn in the corresponding face of $G[\bigcup_{B\in\mathcal{B}}B]$ that contains the facial walk of $G$ between $u_{i}$ and $u_{i+1}$ that is contained in $U$ (note that the vertices of this facial walk are bounded $H$ in $G[\bigcup_{B\in\mathcal{B}}B]$ and the only two other vertices incident to this face are $v_{i},v_{i+1}$). Since $\mathcal{B}$ is a 4-surround of $(G,\mathcal{H})$, it follows that for every $H^{\prime}\in\mathcal{H}$ that contains a vertex incident to $F_{v_{i}v_{i+1}}$, the distance between $H$ and $H^{\prime}$ in $\textbf{RIG}(G,\mathcal{H}\backslash\mathcal{H}_{0})$ is at most $4$. This in particular implies that $F_{v_{i}v_{i+1}}$ will satisfy the 4th bullet in the definition of a $8$-fortification of $(G,\mathcal{H})$. This in particular implies that no vertex incident to $F_{v_{i}v_{i+1}}$ is incident to the outerface of $G$. Then observe that $R_{F,H}$ is a connected set in $G\cup E_{F,H}$. For an illustration of such a $R_{F,H},E_{F,H}$, see Fig.˜8. Let $G^{*}$ be the planar graph obtained from adding these edges from the union of all such $E_{F,H}$. Let $\mathcal{R}$ be the collection of all such $R_{F,H}$. Then every element of $\mathcal{R}$ is a connected set of $G^{*}$.
Рис. 8: The (orange) vertex set $R_{F,H}$ given the (blue) encasing $\mathcal{B}$, the (red) $H\in\mathcal{H}$, and $F\in\mathcal{F}$. The orange vertices are the vertices of $R_{F,H}^{\prime}$. The pink edges are $E_{F,H}$, and the arrows orthogonal to them indicate their sides.
@@
Let $G$ be a finite planar graph and $\mathcal{H}$ a spanning collection of connected vertex sets of $G$. Then there is a $(73936,80)$-impression $(G^{\prime},\mathcal{M})$ of $(G^{\prime},\mathcal{H})$, where $G^{\prime}$ is a fortification of $G$ and $\mathcal{H}$ 88-spans $G^{\prime}$.